OpenStudy (linda3):

Find an equation of the circle that has center (1,-6) and passes through (5,-1) .

OpenStudy (linda3):

I will give medal...

OpenStudy (anonymous):

lol u'll get 2 medals, 2 fans, and 2 fan testimonials @linda3 u gotta know how to bribe the people ;)

OpenStudy (linda3):

lol.... too funny... >.<

OpenStudy (anonymous):

gracias mi amor ;)

OpenStudy (linda3):

no... it's stupid OS.... it's broken...

OpenStudy (linda3):

y'd u delete your comment!?? I sound like i'm talking to myself -_-

OpenStudy (amorfide):

you know the centre, you know what it passes through if we connect the centre and the point it passes through we get the radius to do that we need the distance of the segment. To do this we have a formula d=sqrt((x2-x1)^2+(y2-y1)^2) this gives us our radius squared. x2 x1 y2 y1 are the coordinates from the centre and point it passes through once you get d use your standard equation (x+a)^2+(y+b)^2=r^2 where r^2=d

OpenStudy (linda3):

hmm... I'm looking for a formula that looks more like this.. (x-h) to the powers of 2 + (y-k) to the power of 2 = r to the power of 2

OpenStudy (amorfide):

it is the same formula with different letters lol

OpenStudy (linda3):

oh lol ... ok...

OpenStudy (zubhanwc3):

ok, lets start with the equation \[(x-h)^{2}+(y-k)^{2}=r\]

OpenStudy (zubhanwc3):

with the center being (1,-6) it becomes \[(x-1)^{2}+(y+6)^{2}=r ^{2}\]

OpenStudy (linda3):

now how do we get r to the power of 2 ?

OpenStudy (zubhanwc3):

to find r^2, lets do a distance formula. \[\sqrt{(x _{2}-x _{1}) +(y _{2}-y _{1})}\]

OpenStudy (anonymous):

I wanted to draw it out, but since it's not working, imagine the center at (1, -6) and the other point at (5, -1). The distance between them must be the radius. By the distance formula, r = sqrt( (5-1)^2 + (-1+6)^2 ) = sqrt(4^2 + 5^2) = sqrt(41). By looking at the center, we see that that the circle is shifted 1 to the right (x - 1) and 6 down (y + 6). Plug all this into the equation for a circle that you mentioned and you get. (x-1)^2 + (y+6)^2=r^2 Plug in sqrt(41) for r (which we got earlier) to get the final answer: (x-1)^2 + (y+6)^2=41

OpenStudy (amorfide):

use the formula for the distance from two points

OpenStudy (linda3):

now there's some talking! hahaha

OpenStudy (zubhanwc3):

once u get that answer, u replace r with the answer,

OpenStudy (zubhanwc3):

points are (1,-6) (5,-1) \[\sqrt{(5-1)+(-1-6)}\]

OpenStudy (zubhanwc3):

\[-1--6\]*

OpenStudy (linda3):

i really wish the equation signs on OS would work cause it make this a lot easier .. -_-

OpenStudy (zubhanwc3):

it gives u \[\sqrt{9}\] or 3

OpenStudy (linda3):

wait... but thats a different answer than what @bagajr gave

OpenStudy (zubhanwc3):

so the equation is (x−1)2+(y+6)2=r2\[(x-1)^{2}+(y+6)^{2}=3^{2}\]

OpenStudy (linda3):

so does r to the power of 2 equal 3 or 41?

OpenStudy (zubhanwc3):

weird. the distance formula as i said, was stated above, ima double check my work

OpenStudy (anonymous):

You forgot the squares in distance formula

OpenStudy (zubhanwc3):

no, it equals 3^2 or 9

OpenStudy (zubhanwc3):

o. right i did lol

OpenStudy (anonymous):

So do we agree on r^2=41?

OpenStudy (linda3):

gahhh... which answer is right O_o

OpenStudy (zubhanwc3):

correction on distance formula. \[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (linda3):

i'm gonna go with 41 as the right answer for r to the power of 2

OpenStudy (amorfide):

41=r^2

OpenStudy (amorfide):

sqrt to get r

OpenStudy (zubhanwc3):

bagajr is

OpenStudy (zubhanwc3):

so u can right the answer as \[(x-1)^{2}+(y+6)^{2}=1681\] or \[41\] if u prefer

OpenStudy (zubhanwc3):

write*

OpenStudy (amorfide):

that is wrong

OpenStudy (zubhanwc3):

41^2*

OpenStudy (amorfide):

no square root 41 is r

OpenStudy (zubhanwc3):

no, r = 41.

OpenStudy (amorfide):

5-1=4 -1--6=5 root 41=r 41=r^2

OpenStudy (zubhanwc3):

41 is the distance, and r^2 is that distance squared

OpenStudy (amorfide):

root41 is the distance

OpenStudy (linda3):

no i believe amor is actually right

OpenStudy (amorfide):

I am right

OpenStudy (zubhanwc3):

lol >.< probably

OpenStudy (zubhanwc3):

forgot to square root in distanec formula T_T

OpenStudy (linda3):

lol its okay! xD thank you for helping me!!! :)

OpenStudy (amorfide):

http://answers.yahoo.com/question/index?qid=20110923162212AAMSUSG maybe this will help clear things

OpenStudy (zubhanwc3):

i should do stuff on paper moer often, rather than in my head T_T

OpenStudy (amorfide):

pen and paper is your best friend ;)

OpenStudy (zubhanwc3):

ive become lazy since i started online school :/

OpenStudy (amorfide):

lmao when I am on my computer I find it so much effort to get pen and paper xD

OpenStudy (linda3):

thank you also @amorfide!

OpenStudy (amorfide):

No problem, anything to procrastinate my studies for my exam on friday

OpenStudy (linda3):

haha yeah i have an exam next wednesday! :/

OpenStudy (amorfide):

Guys important question... where is my calculator? It was on my bed, and then it vanished, medal to correct answer

OpenStudy (zubhanwc3):

it fell from the bed, was kicked away, and is now in the kitchen.

OpenStudy (amorfide):

lol it was in the kitchen aswell